# Hypergames: Modeling Misaligned Perceptions and Nested Beliefs for Multi-agent Systems

### v1

@@ 1 Introduction
Strategic decision-making is a key issue in many real-world scenarios involving interactions among multiple agents, and as such is a fundamental agentic capability. Since agents in these settings must choose actions based on the actions of others, game theory naturally emerges as the discipline providing formal tools to model, design, and analyze such systems [58]. However, real-world environments are often marked by uncertainty, misaligned perceptions, and nested beliefs, which frequently render traditional models inadequate. To address these limitations, researchers have developed frameworks that incorporate cognitive constraints, recursive reasoning, and heterogeneous agents exhibiting bounded rationality [33, 63, 54, 15, 20].
Simpler and less expressive models, such as the k-level theory and the cognitive hierarchy model [15], acknowledge differences in players’ reasoning capacities but offer limited insight into the structure and underlying nature of these differences. Epistemic game theory provides a formal language for representing player types, belief hierarchies, and uncertainty; however, its predominantly probabilistic models only partially account for the mechanisms that give rise to these cognitive variations [20].
In contrast, hypergame theory advances the game-theoretic paradigm by explicitly modeling subjective representations of the strategic scenario – so-called perceptual games – in which agents may disagree about the components of the game: the participating players, the available actions, or the payoff structure [7]. This capability makes hypergames particularly suitable for multi-agent system (MAS) frameworks that aim to emulate human-like reasoning and cognitive sophistication. Such suitability aligns with recent interest in theory of mind (ToM) – a psychological model of human reasoning that characterizes the ability to attribute mental states to oneself and others [57] – and its integration into agentic artificial intelligence [59].
@@ 2.1 Game-theoretic Foundations
While decision theory provides tools to reason about how an individual agent can derive choices, when another decision-maker is involved the player may have to explicitly reason about how the opponent is trying to solve the same problem. Game theory is better equipped for such analyses, as it deals with conflicts where decisions and their outcomes are interdependent between agents [53].
@@ 2.1.1 Elements of a Game
[58] introduces the key elements of game-theoretic interactions through the acronym PAPI: Players, Actions, Payoff and Information. We define these four elements as follows:
@@
Players are the decision-making entities involved in the strategic interaction. Game-theoretic models typically assume that players are [53]:
@@
Actions represent the immediate choices available to each player in the game they are involved with. In contrast, strategies are agents’ plans of actions based on the observed interaction history [56]. We distinguish between two classes of strategies [51]:
@@
Payoffs quantify the preferences of each player over possible outcomes. They are typically represented as real-valued utility functions $U_{i}:A\to\mathbb{R}$, where $A$ is the joint action space. Numerical payoffs enable the use of optimization and equilibrium analysis [78, 30].
@@
  • Perfect vs. Imperfect Information: Whether agents can fully observe others’ actions [56];
  • Complete vs. Incomplete Information: Whether agents can fully observe others’ payoffs [30];
  • Symmetric vs. Asymmetric Information: Whether agents can access the same quality and quantity of information [58].
@@ 2.1.2 Formalizing Games
    $u_{i}:A\rightarrow\mathbb{R}\text{ defined as: }$:
@@
When players employ mixed strategies – probability distributions over their pure actions – the outcome of the game becomes probabilistic. The expected utility for a player is then calculated as the average payoff, weighted by the likelihood of each joint action occurring under the players’ chosen strategies [78]. Formally, given mixed strategies $\sigma_{i}$ and $\sigma_{j}$ for players $i$ and $j$, the expected utility for player $i$ is:
@@
[…]
@@
In the case of rock-paper-scissors, no pure strategy profile constitutes a Nash equilibrium. However, a unique mixed strategy Nash equilibrium exists where each player chooses Rock, Paper, and Scissors with equal probability [10].
@@ 2.2 Games of Misperception
Real-world agents may differ in their abilities to observe and process information, potentially leading to distinct internal models of others and misaligned understandings of the conflict, the available actions, or the associated outcomes. Hypergames provide an explicit model of such perceptions by capturing each agent’s subjective interpretation of the situation [8]. Furthermore, hypergame analyses rely on preference relations rather than numerical payoffs, resulting in a qualitative model of desirability. This approach aims to mitigate challenges associated with assigning precise numerical values to real-world outcomes.
Hypergames are considered models of bounded rationality [63, 76] because they support reasoning in which an agent may perceive an opponent’s choice as irrational, even though that choice is rational from the opponent’s own perspective [45]. Formally, a hypergame is composed of each player’s perceptual game – a version of the conflict that represents the player’s interpretation – and is defined as the tuple $H=(N,\{G_{i}\}_{i\in N})$ where:
@@
        For outcomes $x,y\in A_{i}$, $x>y$ indicates that $i$ prefers outcome $x$ to outcome $y$.
@@
This example captures a situation where Scissors is not recognized as an available action by Agent 2, whereas Agent 1 perceives the complete action space. This form of perceptual asymmetry mirrors the classic ”Fall of France” hypergame analysis [6]. The resulting discrepancy in perceived game structure can lead to unexpected outcomes – referred to in hypergame theory as strategic surprise [8].
@@ 2.3 Hierarchical Hypergames
Hypergame theory introduces a framework for modeling individual agents’ subjective interpretations of the conflict they are involved in. By incorporating higher-order perceptions [8], hypergames enable the formalization of recursive reasoning – e.g., a decision-maker $r$ might consider what $q$ believes about how $p$ perceives the situation. These are called $m$-th order expectations, where $m$ corresponds to the depth of the perceptual chain.  [81] later formalized and extended Bennet’s framework into a hierarchical hypergame analysis structure, addressing the overly abstract nature of earlier definitions and proposing a more rigorous, standardized representation of multi-level perceptual games. In this extended formalism, a $0^{th}$-level hypergame is equivalent to a standard game-theoretic without misaligned perceptions and is defined as a tuple $H^{0}=(N,P,S,U,V)$, where:
@@
At each level, the formal structure is extended by an additional layer of subjective (hyper)games built on nested perceptions. While the mathematical framework supports $L^{th}$-order hypergames of arbitrary depth, we restrict the discussion to level 3, reflecting empirical observations that human strategic reasoning seldom exceeds three levels [15] and balancing expressiveness and computational tractability.
@@
The Hypergame Nash Equilibrium (HNE) extends the classical Nash Equilibrium to settings where players may have divergent perceptions of the game [81]. The HNE formalizes mutual best responses in this context of misaligned perceptions; that is, the HNE is an NE in each player’s corresponding subjective game. Given a strategy profile $a^{*}=(a^{*}_{i},a^{*}_{-i})$:
@@ 2.4 Hypergame Normal Form
The normal form is a foundational representation in game theory used to express strategic interactions among decision-makers, where players’ actions and the corresponding payoffs are captured in a table, typically referred to as the payoff matrix [58]. The Hypergame Normal Form (HNF) extends the standard normal form to capture a single player’s beliefs – drawing on decision-theoretic concepts [32] – about their opponent’s reasoning [77]. Originally proposed for analyzing complex military planning scenarios, HNF enables an agent to reason over multiple possible opponent strategies, each corresponding to a different perceptual subgame. Table 1 illustrates an HNF matrix for the rock-paper-scissors game used in previous examples from the row player’s perspective. The HNF consists of:
@@
The game in Table 1 defines the row player’s beliefs about the column player in rock-paper-scissors. $C_{0}$ corresponds to the mixed strategy Nash equilibrium (uniform randomization), while $C_{1}$ reflects empirical distributions derived from online gameplay data [10]. In $C_{2}$, the row player believes that the opponent cannot play Scissors. The belief context assigns the highest probability to $C_{2}$, indicating that the row player expects this restricted subgame to be most likely. The RMS entries mirror responses to each $C_{k}$. Additionally, the HNF matrix can include expected utility (EU) and hypergame expected utility (HEU) calculations as additional rows beneath the matrix – omitted in Table 1 for clarity.
@@
Hyperstrategies are mixed strategies over RMSs, represented as vectors of assigned probabilities adjacent to the RMS quadrant. We include a hypothetical hyperstrategy $H_{1}$ in Table 1 for illustration. Vane [77] defines four hyperstrategies that meet the following efficiency criterion:
@@
Vane [75] introduced hypergame expected utility (HEU) as a bridge between decision theory and game theory. HEU incorporates a parameter $g$ – the “fear of being outguessed” – and the worst-case outcome $G$, defined as the lowest expected utility for the row player. Then, given a hyperstrategy $hs$, we define the HEU as:
@@
The relative performance of hyperstrategies such as MO, PS, and WS varies with the uncertainty parameter $g$, with NEMS becoming increasingly dominant as uncertainty rises. The analysis of $g$’s effects – detailed in [75] – suggests that when agents hold accurate beliefs about their opponent’s strategy, HEU-optimized hyperstrategies can outperform standard equilibrium analysis.
@@ 3 Hypergame Theory in Multi-Agent Systems
While hypergame theory is a model of bounded rationality that can provide multi-agent systems with a framework that can model agent heterogeneity and environmental uncertainties to a greater level of detail than probabilistic methods, it was originally designed to be an analytical framework targeting post-hoc studies of complex conflicts instead of providing tools for the decision-maker on the spot. While hypergame models have yet to become prevalent in dynamic settings multi-agent systems entail, they have a potential to overcome the constraints posed by typical game-theoretic models on the complexity agent-based frameworks can capture [61]. Motivated by this identified potential, in this section we provide a systematic review of hypergame-theoretic works focusing on studies that entail agent-based applications or the potential to be integrated into dynamic agentic systems.
@@ 3.1 Methodology
We removed 303030 entries due to duplicity. Removed entries included different versions of the same manuscript (e.g., preprints and published versions) as well as overlapping content such as thesis chapters and standalone publications.
@@ 3.2 Hypergame-theoretic Publication Trends
Interestingly, [45] stated that “Unfortunately very little in the way of hypergame theory application has been done in the arena of cyber warfare”, while a decade later the current shift is dominantly lead by attack-defense modeling and simulations in cybersecurity – as demonstrated by Figures 679 and the various cybersecurity-related surveys covering hypergame-based approaches we discuss in section 5.
@@ 3.3 Taxonomy of Hypergame Models
We also distinguished works by whether they were inspired by the original [8] or extended [81] hypergame framework – collectively denoted as Multi-level Hypergames (MLH) – or based on the Hypergame Normal Form (HNF). We found 7 papers that refer to HNF, 35 rely on Multi-level Hypergames, and 2 papers use components from both concepts, which we denote as Hybrid approaches.
@@ 3.4.1 Application Domains
Game Theory Works that advance game theory in general, without applying to any specific practical domain, are categorized under the general domain of Game Theory. These papers utilize hypergame-theoretic tools to define equilibrium concepts, model learning dynamics under misperception, or devise strategic responses to asymmetry and deception. For example,  [64] integrates Fuzzy Game Theory with hypergames by incorporating fuzzy strategies and imprecise preference perceptions. In contrast,  [83] focuses on Deception and uses hypergames to analyze how misinformation affects equilibrium strategies. Additionally,  [52] proposes a hypergame-based Q-learning approach to solve Pursuit-Evasion games, while  [62] develops a hypergame-theoretic model of deception in concurrent Stochastic Games. Lastly, various works leverage the hypergame-theoretic Evolution of Beliefs in adversarial games characterized by misperceptions. These studies define mechanisms for belief updates [24, 28] that target the learning of equilibria [25, 26] and belief manipulation [27].
@@
Cybersecurity Cybersecurity represents the most active domain for agent-compatible hypergame applications. This prevalence can be attributed to the inherent suitability of hypergames for capturing misinformation and modeling active deception – both critical issues within cybersecurity contexts. The majority of the surveyed works address Attack-Defense Modeling in general, formulating hypergames on graphs and applying linear temporal logic to guide deceptive strategy synthesis [41, 42, 47, 50]. Other examples include belief manipulation [29, 4], mission impact assessments employing hypergame-theoretic frameworks [73, 84], and other specialized hypergame-based methodologies for attack-defense modeling [34, 23, 5, 14]. In addition, several works target a narrower area of attack-defense modeling. Defensive Deception entails the use of deceptive strategies to mislead the attacker, often focusing on optimizing strategic honeypot, decoy, and false target assignment [82, 43, 79, 2, 39, 66]. Alternative approaches include using defensive deception strategies to increase the cost of attacks through modeling uncertainty via hypergames [37] and maximizing strategic surprise via sensor revealing strategies [74], or exploring two-way misinformation in deception games [17] and defensive deception against multiple attackers [80].
@@
Social Simulations Social simulations involve multi-agent-based modeling of strategic interactions within conceptualized social environments. The framework proposed by [69] employs a hypergame-based Theory of Mind reasoning framework. This approach has been utilized by [70] to explore the ability of agentic large language models (LLMs) to engage in human-like reasoning during guessing games. Furthermore, [46] leverages hypergame theory to model Coalition Detection in natural language reasoning tasks performed by LLMs.
@@
Robotics In the realm of robotic Control Theory, hypergame applications, for example, focus on modeling obstructed perceptions in traffic scenarios using subjective games [48]. In contrast, [40] aim to develop hypergame-based strategies in asymmetrical information environments between robots and their surroundings. Other studies in robotics explore the implementation of hypergames in the context of Autonomous Vehicles, particularly in the design of deceptive defense strategies for unmanned ground vehicles that operate under asymmetric perception conditions [49]. Others address the cognitive alignment of unmanned aerial vehicles during defensive [35, 36] or exploratory [18] missions.
@@
Multi-agent Systems We categorize applications related to general issues within MAS under this domain. [1] defines a set of hidden-role games targeting deceptive objectives and equipping agents with second-order Theory of Mind conceptualized as second-order hypergames. This framework allows agents to act based on misaligned perceptions of their roles and objectives, facilitating the emergence of asymmetric beliefs and belief manipulation within partially observable, mixed cooperative-competitive settings.
@@
Communications Within the communications sector, we examine decentralized multi-user wireless networks composed of transmitters and receivers whose interactions are limited by bandwidth, channel quality, and computational resources. Semantic communication described a paradigm where only the most task-relevant information is transmitted while receivers utilize reasoning and reconstruction techniques to infer any missing data. In this context, [71] applies hypergame theory to model and manage perceptual discrepancies among agents explicitly.

### v2

# A Survey on Hypergame Theory: Modeling Misaligned Perceptions and Nested Beliefs for Multi-agent Systems

@@ 1 Introduction
Strategic decision-making is a key issue in many real-world scenarios involving interactions among multiple agents, and as such is a fundamental agentic capability. Since agents in these settings must choose actions based on the actions of others, game theory naturally emerges as the discipline providing formal tools to model, design, and analyze such systems [Rasmusen1990-RASGAI]. However, real-world environments are often marked by uncertainty, misaligned perceptions, and nested beliefs, which frequently render traditional models inadequate. To address these limitations, researchers have developed frameworks that incorporate cognitive constraints, recursive reasoning, and heterogeneous agents exhibiting bounded rationality [Harsanyi1967, simon1972, nagel95, camerer2004cognitivehieararchy, DEKEL2015epistemicgametheory].
Simpler and less expressive models, such as the k-level theory and the cognitive hierarchy model [camerer2004cognitivehieararchy], acknowledge differences in players’ reasoning capacities but offer limited insight into the structure and underlying nature of these differences. Epistemic game theory provides a formal language for representing player types, belief hierarchies, and uncertainty; however, its predominantly probabilistic models only partially account for the mechanisms that give rise to these cognitive variations [DEKEL2015epistemicgametheory].
In contrast, hypergame theory advances the game-theoretic paradigm by explicitly modeling subjective representations of the strategic scenario – so-called perceptual games – in which agents may disagree about the components of the game: the participating players, the available actions, or the payoff structure [Bennett1977]. This capability makes hypergames particularly suitable for multi-agent system (MAS) frameworks that aim to emulate human-like reasoning and cognitive sophistication. Such suitability aligns with recent interest in theory of mind (ToM) – a psychological model of human reasoning that characterizes the ability to attribute mental states to oneself and others [Premack_Woodruff_1978theoryofmind] – and its integration into agentic artificial intelligence [rocha2023tomReview].
@@ 2.1 Game-theoretic Foundations
While decision theory provides tools to reason about how an individual agent can derive choices, when another decision-maker is involved the player may have to explicitly reason about how the opponent is trying to solve the same problem. Game theory is better equipped for such analyses, as it deals with conflicts where decisions and their outcomes are interdependent between agents [myerson1984introduction].
@@ 2.1.1 Elements of a Game
[Rasmusen1990-RASGAI] introduces the key elements of game-theoretic interactions through the acronym PAPI: Players, Actions, Payoff and Information. We define these four elements as follows:
@@
Players are the decision-making entities involved in the strategic interaction. Game-theoretic models typically assume that players are [myerson1984introduction]:
@@
Actions represent the immediate choices available to each player in the game they are involved with. In contrast, strategies are agents’ plans of actions based on the observed interaction history [OsborneRubinstein1994]. We distinguish between two classes of strategies [LuceRaiffa1989games]:
@@
Payoffs quantify the preferences of each player over possible outcomes. They are typically represented as real-valued utility functions $U_{i}:A\to\mathbb{R}$, where $A$ is the joint action space. Numerical payoffs enable the use of optimization and equilibrium analysis [vonNeumann1944, Gibbons1992].
@@
  • Perfect vs. Imperfect Information: Whether agents can fully observe others’ actions [OsborneRubinstein1994];
  • Complete vs. Incomplete Information: Whether agents can fully observe others’ payoffs [Gibbons1992];
  • Symmetric vs. Asymmetric Information: Whether agents can access the same quality and quantity of information [Rasmusen1990-RASGAI].
@@ 2.1.2 Formalizing Games
    $u_{i}:A\rightarrow\mathbb{R}\text{ defined as: }$
@@
When players employ mixed strategies – probability distributions over their pure actions – the outcome of the game becomes probabilistic. The expected utility for a player is then calculated as the average payoff, weighted by the likelihood of each joint action occurring under the players’ chosen strategies [vonNeumann1944]. Formally, given mixed strategies $\sigma_{i}$ and $\sigma_{j}$ for players $i$ and $j$, the expected utility for player $i$ is:
@@
The Nash Equilibrium (NE) is a fundamental solution concept in game theory, capturing the state when neither player can improve their payoff by a unilateral change of strategy [Nash1950equi]. Formally, a strategy profile $a^{*}=(a^{*}_{i},a^{*}_{-i})$ – with i-i denoting the opponent(s) – is a NE, for each player $i\in N$:
@@
In the case of rock-paper-scissors, no pure strategy profile constitutes a Nash equilibrium. However, a unique mixed strategy Nash equilibrium exists where each player chooses Rock, Paper, and Scissors with equal probability [batzilis19rockpaperscissors].
@@ 2.2 Games of Misperception
Real-world agents may differ in their abilities to observe and process information, potentially leading to distinct internal models of others and misaligned understandings of the conflict, the available actions, or the associated outcomes. Hypergames provide an explicit model of such perceptions by capturing each agent’s subjective interpretation of the situation [Bennett1980]. Furthermore, hypergame analyses rely on preference relations rather than numerical payoffs, resulting in a qualitative model of desirability. This approach aims to mitigate challenges associated with assigning precise numerical values to real-world outcomes.
Hypergames are considered models of bounded rationality [simon1972, vane2006advances] because they support reasoning in which an agent may perceive an opponent’s choice as irrational, even though that choice is rational from the opponent’s own perspective [kovach2015hypergamereview]. Formally, a hypergame is composed of each player’s perceptual game – a version of the conflict that represents the player’s interpretation – and is defined as the tuple $H=(N,\{G_{i}\}_{i\in N})$ where:
@@
        For outcomes $\in A_{i}, x>yx>y$ indicates that $i$ prefers outcome $x$ to outcome $y$.
@@
This example captures a situation where Scissors is not recognized as an available action by Agent 2, whereas Agent 1 perceives the complete action space. This form of perceptual asymmetry mirrors the classic ”Fall of France” hypergame analysis [BennettDando1979]. The resulting discrepancy in perceived game structure can lead to unexpected outcomes – referred to in hypergame theory as strategic surprise [Bennett1980].
@@ 2.3 Hierarchical Hypergames
Hypergame theory introduces a framework for modeling individual agents’ subjective interpretations of the conflict they are involved in. By incorporating higher-order perceptions [Bennett1980], hypergames enable the formalization of recursive reasoning – e.g., a decision-maker $r$ might consider what $q$ believes about how $p$ perceives the situation. These are called $m$-th order expectations, where $m$ corresponds to the depth of the perceptual chain.  [Wang1988] later formalized and extended Bennet’s framework into a hierarchical hypergame analysis structure, addressing the overly abstract nature of earlier definitions and proposing a more rigorous, standardized representation of multi-level perceptual games. In this extended formalism, a $0^{th}$-level hypergame is equivalent to a standard game-theoretic without misaligned perceptions and is defined as a tuple $H^{0}=(N,P,S,U,V)$, where:
@@
At each level, the formal structure is extended by an additional layer of subjective (hyper)games built on nested perceptions. While the mathematical framework supports $L^{th}$-order hypergames of arbitrary depth, we restrict the discussion to level 3, reflecting empirical observations that human strategic reasoning seldom exceeds three levels [camerer2004cognitivehieararchy] and balancing expressiveness and computational tractability.
@@
The Hypergame Nash Equilibrium (HNE) extends the classical Nash Equilibrium to settings where players may have divergent perceptions of the game [Wang1988]. The HNE formalizes mutual best responses in this context of misaligned perceptions; that is, the HNE is an NE in each player’s corresponding subjective game. Given a strategy profile $a^{*}=(a^{*}_{i},a^{*}_{-i})$:
@@ 2.4 Hypergame Normal Form
The normal form is a foundational representation in game theory used to express strategic interactions among decision-makers, where players’ actions and the corresponding payoffs are captured in a table, typically referred to as the payoff matrix [Rasmusen1990-RASGAI]. The Hypergame Normal Form (HNF) extends the standard normal form to capture a single player’s beliefs – drawing on decision-theoretic concepts [Hansson2011] – about their opponent’s reasoning [vane2000thesis]. Originally proposed for analyzing complex military planning scenarios, HNF enables an agent to reason over multiple possible opponent strategies, each corresponding to a different perceptual subgame. Table 1 illustrates an HNF matrix for the rock-paper-scissors game used in previous examples from the row player’s perspective. The HNF consists of:
@@
The game in Table 1 defines the row player’s beliefs about the column player in rock-paper-scissors. $C_{0}$ corresponds to the mixed strategy Nash equilibrium (uniform randomization), while $C_{1}$ reflects empirical distributions derived from online gameplay data [batzilis19rockpaperscissors]. In $C_{2}$, the row player believes that the opponent cannot play Scissors. The belief context assigns the highest probability to $C_{2}$, indicating that the row player expects this restricted subgame to be most likely. The RMS entries mirror responses to each $C_{k}$. Additionally, the HNF matrix can include expected utility (EU) and hypergame expected utility (HEU) calculations as additional rows beneath the matrix – omitted in Table 1 for clarity.
@@
Hyperstrategies are mixed strategies over RMSs, represented as vectors of assigned probabilities adjacent to the RMS quadrant. We include a hypothetical hyperstrategy $H_{1}$ in Table 1 for illustration. Vane [vane2000thesis] defines four hyperstrategies that meet the following efficiency criterion:
@@
Vane [vane2000dtgt] introduced hypergame expected utility (HEU) as a bridge between decision theory and game theory. HEU incorporates a parameter $g$ – the “fear of being outguessed” – and the worst-case outcome $G$, defined as the lowest expected utility for the row player. Then, given a hyperstrategy $hs$, we define the HEU as:
@@
The relative performance of hyperstrategies such as MO, PS, and WS varies with the uncertainty parameter $g$, with NEMS becoming increasingly dominant as uncertainty rises. The analysis of $g$’s effects – detailed in [vane2000dtgt] – suggests that when agents hold accurate beliefs about their opponent’s strategy, HEU-optimized hyperstrategies can outperform standard equilibrium analysis.
@@ 3 Hypergame Theory in Multi-Agent Systems
While hypergame theory is a model of bounded rationality that can provide multi-agent systems with a framework that can model agent heterogeneity and environmental uncertainties to a greater level of detail than probabilistic methods, it was originally designed to be an analytical framework targeting post-hoc studies of complex conflicts instead of providing tools for the decision-maker on the spot. While hypergame models have yet to become prevalent in dynamic settings multi-agent systems entail, they have a potential to overcome the constraints posed by typical game-theoretic models on the complexity agent-based frameworks can capture [Sasaki2021multiagaentdecisionsystems]. Motivated by this identified potential, in this section we provide a systematic review of hypergame-theoretic works focusing on studies that entail agent-based applications or the potential to be integrated into dynamic agentic systems.
@@ 3.1 Methodology
We removed $30$ entries due to duplicity. Removed entries included different versions of the same manuscript (e.g., preprints and published versions) as well as overlapping content such as thesis chapters and standalone publications.
@@ 3.2 Hypergame-theoretic Publication Trends
Interestingly, [kovach2015hypergamereview] stated that “Unfortunately very little in the way of hypergame theory application has been done in the arena of cyber warfare”, while a decade later the current shift is dominantly lead by attack-defense modeling and simulations in cybersecurity – as demonstrated by Figures 679 and the various cybersecurity-related surveys covering hypergame-based approaches we discuss in section 5.
@@ 3.3 Taxonomy of Hypergame Models
We also distinguished works by whether they were inspired by the original [Bennett1980] or extended [Wang1988] hypergame framework – collectively denoted as Multi-level Hypergames (MLH) – or based on the Hypergame Normal Form (HNF). We found 7 papers that refer to HNF, 35 rely on Multi-level Hypergames, and 2 papers use components from both concepts, which we denote as Hybrid approaches.
@@ 3.4.1 Application Domains
Game Theory Works that advance game theory in general, without applying to any specific practical domain, are categorized under the general domain of Game Theory. These papers utilize hypergame-theoretic tools to define equilibrium concepts, model learning dynamics under misperception, or devise strategic responses to asymmetry and deception. For example,  [song2009fuzzyhypergame] integrates Fuzzy Game Theory with hypergames by incorporating fuzzy strategies and imprecise preference perceptions. In contrast,  [yuan2024equilibrium] focuses on Deception and uses hypergames to analyze how misinformation affects equilibrium strategies. Additionally,  [li2024hypergameQlearning] proposes a hypergame-based Q-learning approach to solve Pursuit-Evasion games, while  [chongyang2023stochasticgames] develops a hypergame-theoretic model of deception in concurrent Stochastic Games. Lastly, various works leverage the hypergame-theoretic Evolution of Beliefs in adversarial games characterized by misperceptions. These studies define mechanisms for belief updates [gharesifard2010evolution, gharesifard2012evolution] that target the learning of equilibria [gharesifard2011exploration, gharesifard2011learningequilibria] and belief manipulation [gharesifard2011stealthy].
@@
Cybersecurity Cybersecurity represents the most active domain for agent-compatible hypergame applications. This prevalence can be attributed to the inherent suitability of hypergames for capturing misinformation and modeling active deception – both critical issues within cybersecurity contexts. The majority of the surveyed works address Attack-Defense Modeling in general, formulating hypergames on graphs and applying linear temporal logic to guide deceptive strategy synthesis [kulkarni2021reachabilitygames, kulkarni2021hypergamegraphs, kulkarni2021labeling, lening2023dynamichypergames]. Other examples include belief manipulation [gharesifard2014stealthydeception, bakker2020infomanipulation], mission impact assessments employing hypergame-theoretic frameworks [thukkaraju2023mia, yoon2025mia], and other specialized hypergame-based methodologies for attack-defense modeling [house2010hypergame, guiterrez2018, bakker2021metagames, cheng2022stackelberg]. In addition, several works target a narrower area of attack-defense modeling. Defensive Deception entails the use of deceptive strategies to mislead the attacker, often focusing on optimizing strategic honeypot, decoy, and false target assignment [bowei2020iobt, kulkarni2020decoyallocation, wan2021foureyedefensivedeceptionbased, anwar2020, kulkarni2024integratedresourceallocationstrategy, shen2025iot]. Alternative approaches include using defensive deception strategies to increase the cost of attacks through modeling uncertainty via hypergames [jia2024defense] and maximizing strategic surprise via sensor revealing strategies [udupa2024reactivesynthesissensorrevealing], or exploring two-way misinformation in deception games [Cho2019deceptiongames] and defensive deception against multiple attackers [wan2023APT].
@@
Social Simulations Social simulations involve multi-agent-based modeling of strategic interactions within conceptualized social environments. The framework proposed by [trencsenyi2025approximatinghumanstrategicreasoning] employs a hypergame-based Theory of Mind reasoning framework. This approach has been utilized by [trencsenyi2025influencehumaninspiredagenticsophistication] to explore the ability of agentic large language models (LLMs) to engage in human-like reasoning during guessing games. Furthermore, [kulkarni2025coalitiondetection] leverages hypergame theory to model Coalition Detection in natural language reasoning tasks performed by LLMs.
@@
Robotics In the realm of robotic Control Theory, hypergame applications, for example, focus on modeling obstructed perceptions in traffic scenarios using subjective games [kahn2022traffic]. In contrast, [kulkarni2019reactivegames] aim to develop hypergame-based strategies in asymmetrical information environments between robots and their surroundings. Other studies in robotics explore the implementation of hypergames in the context of Autonomous Vehicles, particularly in the design of deceptive defense strategies for unmanned ground vehicles that operate under asymmetric perception conditions [lv2024synthesisforUAV]. Others address the cognitive alignment of unmanned aerial vehicles during defensive [he2022leader, hai2023uavcommunication] or exploratory [dharmadhikari2021dronesexploring] missions.
@@
Multi-agent Systems We categorize applications related to general issues within MAS under this domain. [aitchison2021hiddenrolegames] defines a set of hidden-role games targeting deceptive objectives and equipping agents with second-order Theory of Mind conceptualized as second-order hypergames. This framework allows agents to act based on misaligned perceptions of their roles and objectives, facilitating the emergence of asymmetric beliefs and belief manipulation within partially observable, mixed cooperative-competitive settings.
@@
Communications Within the communications sector, we examine decentralized multi-user wireless networks composed of transmitters and receivers whose interactions are limited by bandwidth, channel quality, and computational resources. Semantic communication described a paradigm where only the most task-relevant information is transmitted while receivers utilize reasoning and reconstruction techniques to infer any missing data. In this context, [thomas2024hypergametheorydecentralizedresource] applies hypergame theory to model and manage perceptual discrepancies among agents explicitly.